OpenStudy (anonymous):

choose the standard form of the equation y=-1/4x-6

OpenStudy (anonymous):

Standard form of a line is \[Ax + By = C\]\[y = -\frac{1}{4}x - 6 \\ y + \frac{1}{4}x = -6 \\ \frac{1}{4}x + y = -6\]